# Energetic particles accelerated via turbulent magnetic reconnection in protoplanetary discs – I: ionisation rates

### v1

@@ 2.2.2 Turbulence modelling in a static disc model
To evaluate the energy distribution of non-thermal particles, we first summarise the key properties of HD/MHD turbulence in the region of interest, focusing on MRI-driven turbulence. MRI requires sufficient ionisation for strong gas–field coupling. The geometric setup adopted here is sketched on the left part of Fig. 1.
@@
The velocity amplitude depends both on $\alpha_{\rm eff}$ and on how this efficiency partitions between spatial and velocity scales, as discussed in Hughes et al. (2011). If eddies act on scales $\sim H$, one expects $varv_{\rm turb}\sim\alpha_{\rm eff}c_{s}$; if $\alpha_{\rm eff}$ distributes similarly across length and velocity, then $varv_{\rm turb}\sim\sqrt{\alpha_{\rm eff}}c_{s}$. Empirically, observations and simulations rather support,
@@
We evaluate $\alpha_{\rm eff}$ using the non-ideal, MRI-based prescription of Thi et al. (2019), described in Appendix A. This yields $varv_{\rm turb}\approx 0.1\text{–}0.4c_{s}$ in the upper layers and $varv_{\rm turb}\lesssim 0.01c_{s}$ near the mid-plane, consistent with observational constraints, and provides the values of the $L$ parameter encoding the reconnecting current sheet length and the turbulent speed used below in section 2.3.2 to connect turbulence to reconnection geometry and, ultimately, to the non-thermal particle distribution.
@@ 2.3.1 Magnetohydrodynamic turbulence and magnetic reconnection
Turbulence by itself does not guarantee magnetic reconnection. However, the atmospheric layers of protoplanetary discs combine physical conditions that make reconnection a natural outcome of the dynamics. Indeed, in this region, irradiation by stellar far UV and X-rays maintains high ionisation fractions ($\gtrsim 0.1$, see the H+/H transition Fig. 1) ensuring tight magnetic coupling (weak Ohmic/ambipolar slippage). Where coupling is good, the MRI efficiently sustains turbulence and Maxwell stresses. The canonical expectation is that MRI is most efficient for high $\beta_{p}$ (¿100); MRI in low-$\beta_{p}$ require additional conditions to grow. The parameter $\beta_{p}=P_{\text{gas}}/P_{\text{mag}}$ is the ratio of the gas pressure to the magnetic pressure. In particular, Kim and Ostriker (2000) show that reducing the toroidal component of the magnetic field re-opens the MRI window even at low $\beta_{p}$. Recent simulations of disc surface layers are consistent with this picture. Under conditions of strong ionisation and suitable field geometry, MRI-driven turbulence naturally generates magnetic reversals, shears and then thin current sheets, the precursors of reconnection (e.g. Jacquemin-Ide et al., 2021; Rosenberg and Ebrahimi, 2021), making this process frequent and energetically relevant in disc atmospheres.
@@
There are only few works linking turbulence and magnetic reconnection in accretion discs. Kadowaki et al. (2018) explore the connection between turbulence and fast magnetic reconnection in accreting systems. The authors conduct 3D ideal MHD simulations in the shearing box approximation. Although ideal, the simulations allow magnetic reconnection thanks to numerical resistivity. Their setup starts from an initial azimuthal magnetic field and an exponential profile of the gas density over a scale $H$. The azimuthal magnetic field component dominates the initial radial and vertical components. The initial plasma beta parameter, $\beta_{p}$, is set to 1, 10 and 100 in a series of three different runs. The simulations follow the evolution of the magnetic components under the effect of the MRI but also of the Parker-Rayleigh-Taylor instability (PRTI, Parker, 1966).
@@
PRTI together with MRI, sustain an $\alpha-\Omega$ dynamo that reaches a quasi-steady state in which the characteristic reconnection speed is $V_{r}\approx(0.13\pm 0.09)V_{A}$ in the disc and $(0.17\pm 0.10)V_{A}$ in the corona, and the typical current-sheet width is $\Delta_{r}\approx 0.08H$ (disc) and 0.10H0.10\,H (corona), where $V_{A}$ is the local Alfvén speed. We compute the scale height using the Keplerian thin disc relation $H=c_{s}/\Omega_{K}$.  [^3] Although the ProDiMo structure is moderately flared ($H/R\!\sim\!0.1$), the deviation of the angular frequency from Keplerian due to pressure support is $\mathcal{O}[(H/R)^{2}]\!\sim\!10^{-2}$ (Armitage, 2019). This percent-level correction is well within typical astrophysical uncertainties, so the Keplerian approximation is adequate for our purposes.
@@
The values of $V_{r},\Delta_{r},H,V_{A},M_{A}$} computed from ProDiMo thus provide a coherent, simulation-supported model from which we estimate the reconnection–driven EP acceleration parameters, as detailed in Sect. 3.2.
@@ 2.3.2 Multi-scale magnetic reconnection modelling and particle acceleration
At scales larger than kinetic ones, magnetic reconnection is allowed through the breaking of magnetic frozen-in conditions by collisions, even when the plasma is fully ionised and neutrals are not present, see e.g. (Sweet, 1958; Furth et al., 1963). While at kinetic scales magnetic reconnection is known to be sustained by electron inertia or the electron pressure tensor and to be efficient (i.e., fast) in converting magnetic energy into particle acceleration, fluid models have struggled to provide an explanation for fast energy release. Different models have been proposed in recent years (see, e.g., the discussion in Lazarian et al. 2020) among which the tearing instability on thin current sheets is the most promising (see e.g. Loureiro et al. 2007; Bhattacharjee et al. 2009; Pucci and Velli 2014). The survival of laminar current sheets, necessary for the onset of the tearing instability, remains controversial in the context of turbulent protoplanetary disks, and outside the scope of this work. Still, reconnection is observed in 2D and also 3D turbulent simulations even with sufficiently high magnetic Reynolds number (see Dong et al. 2022 and references therein). In the latter work, the formation of reconnection-produced magnetic flux ropes in 3D is shown to exhibit a complex morphology, differing significantly from the simpler island-like morphology in 2D.
The authors note the similar size of the flux ropes between the 2D and 3D configurations, though the size of plasmoids is distributed over a wide variety of scales.
@@
The length $L_{r}$ of the largest current sheet forming in a turbulent region is assumed to be approximately equal to the local turbulent injection scale, i.e. $L\approx L_{r}$; largest current sheets would be destroyed by the turbulence. The thickness $\Delta_{r}$ is set by the scale over which a magnetic field line deviates from its original direction (see Lazarian and Vishniac 1999) or, adopting a laminar tearing model, the intrinsic scale at which reconnection becomes fast (Bhattacharjee et al., 2009; Loureiro et al., 2007; Pucci and Velli, 2014). Fig. 2 schematically illustrates our working picture: sub-Alfvénic MRI turbulence ($M_{A}<1$) at the injection scale $L$ (left) generates current sheets where turbulent magnetic reconnection operates (right). The reconnecting current sheet has a typical length $L_{r}$ and thickness $\Delta_{r}$, with $L_{r}=L$.
@@
Figure 2: Sketch of MRI-driven turbulence and embedded turbulent magnetic reconnection. Left: eddies at the injection scale $L$ in a sub-Alfvénic regime characterised by the Alfvénic Mach number $M_{A}$. Right: the reconnection layer produced by the turbulence, with length $L_{r}=L$ and thickness $\Delta_{r}$. The turbulence parameters $L,M_{A}$} set the reconnection, geometry parameters $L_{r},\Delta_{r}$}, speed $V_{\rm rec}\sim V_{A}M_{A}^{2}$ and the EP acceleration parameters used in this work.
@@
In a generic magnetic reconnection geometry, EP acceleration is characterized by three parameters, the acceleration rate $\alpha_{\rm acc}$, the escape time $\tau_{\rm esc}$, and the injection time $\tau_{\rm inj}$, which depend on the reconnection model and on the characteristic size and lifetime of the reconnecting current sheet (Lazarian and Vishniac, 1999; Kowal et al., 2012; Lazarian et al., 2020; Xu and Lazarian, 2023),
@@
For the purpose of this paper we consider the developments of the turbulent-reconnection framework of Zhang et al. (2023) to be appropriate for the estimate of the reconnection parameters. Thus, the turbulent parameters $L,M_{A}$} set the reconnection speed $V_{\rm rec}\sim V_{A}M_{A}^{2}$ (Lazarian and Vishniac, 1999) and, as shown below, the parameters of the EP acceleration model. The particles are considered to be accelerated multiple times within the turbulent layer, transitioning between different current sheets and undergo Fermi acceleration through island compression.
The same assumption holds, assuming the plasmoid instability or any fast tearing instability kicks in. In particular we consider here the “ideal” tearing instability, see Pucci and Velli (2014). Then the acceleration rate is $\alpha_{\rm acc}\sim t^{-1}_{A}$ where the proportionality constant depend on the topology of the reconnecting magnetic field (Pucci et al., 2018) and other non ideal effects, see e.g. Pucci et al. (2017); Shi et al. (2019). In this work, we adopt the results of Xu and Lazarian (2023), giving
@@
In conclusion, comparing collisional fluid modelling to Zhang et al. (2023) brings correction factors to the energy effectively converted within the turbulent layer in the disc’s atmosphere. The development of detailed particle acceleration models based on the interplay of reconnection and turbulence remains largely phenomenological, and various types of dimensional analysis may be adopted. Not knowing the details of the microscopic acceleration process at this stage, we introduce the efficiency parameter $\kappa$ and defer a full treatment to future work.
@@ 3.1 Energy distribution from acceleration by a Fermi-like process
Magnetic reconnection can efficiently accelerate a fraction of the thermal particle population to supra-thermal energies through Fermi-like acceleration mechanism. Assuming such an acceleration process is active in the region of interest, the energy gain by a particle is proportional to the particle kinetic energy, $\dot{E}=\alpha_{\rm acc}E$, with $\alpha_{\rm acc}$ being the acceleration rate. The equation for the energy distribution is (Guo et al., 2014, 2015; Oka et al., 2023),
@@
where $F_{\rm nt}$ is the particle energy distribution (erg-1 cm-3), $\dot{E}$ the particle energy gain rate (erg s-1), $\tau_{\rm inj}$ is the timescale for injection of the initial particle distribution $F_{\rm inj}$ into the accelerating region, and $\tau_{\rm esc}$ is the escape timescale from the acceleration region.
Assuming that the initial upstream distribution is a Maxwellian distribution  [^4] The analysis performed in Guo et al. (2014) has been developed in the context of plasmas at high magnetization, but the general form of the solution does not depend on the magnetization regime, it simply assumes Fermi-I type acceleration, an injection and an escape rate. The solution in Eq. 8 assumes non-relativistic temperature of the background plasma, which is the case in the astrophysical system under consideration.,
@@
where $F_{\rm inj}$ is the upstream particle energy distribution in units of thermal energy, $\epsilon$ is the energy in units of thermal energy, $\epsilon=E/E_{\rm th}$, with $E_{\rm th}=3/2k_{B}T$, where $k_{B}$ is the Boltzmann constant and $T$ the temperature of the medium and $n_{\rm inj}$ is the density of particles entering in the acceleration region.
The solution of Eq. (6) is a non-thermal distribution that can be written in units of thermal energy as,
@@
In the next sections, we will evaluate the acceleration parameters assuming a specific turbulent magnetic reconnection model and see how to normalize it based on the power released by the turbulence in the disc.
@@ 3.2 Energy distribution parameters in turbulent magnetic reconnection
The shape of the non-thermal energy distribution is set by the three parameters, $\alpha_{\rm acc}$, $\tau_{\rm esc}$, and $\tau_{\rm inj}$ as defined in Eq. (5), which depend on the characteristic size and lifetime of the reconnecting current sheet.
@@
After a time $\tau_{\rm CS}$, a power-law component emerges if the factor $\Gamma_{3/2+\beta}\left(\epsilon e^{-\alpha_{\rm acc}\tau_{\rm CS}}\right)-\Gamma_{3/2+\beta}(\epsilon)$ in Eq. (8) is (approximately) energy-independent. This holds across the supra-thermal range provided $\epsilon e^{-\alpha_{\rm acc}\tau_{\rm CS}}\lesssim 1$. For $\epsilon e^{-\alpha_{\rm acc}\tau_{\rm CS}}\gtrsim 1$, the spectrum bends away from a pure power law, defining a finite-time cut-off
@@
which directly reflects Fermi-like energisation with $\dot{\epsilon}=\alpha_{\rm acc}\epsilon$ acting for a duration $\tau_{\rm CS}$.
@@
At the low-energy end, we define the injection threshold $\epsilon_{\rm inj}$ by continuity between thermal and non-thermal components,
@@
With the spectral shape fixed by $\alpha_{\rm acc}$, $\tau_{\rm inj}$, and $\tau_{\rm esc}$ (thus setting $p$, $\epsilon_{\rm inj}$, and $\epsilon_{\max}$), the remaining task is to determine the normalisation, i.e. the number density of non-thermal particles, from the local energy budget. In the next subsection we link the particle power to a fraction of the turbulent accretion dissipation to obtain the non-thermal EP density.

### v2

@@ 2.2.2 Turbulence modelling in a static disc model
To evaluate the energy distribution of non-thermal particles, we first summarise the key properties of MHD turbulence in the region of interest, focusing on MRI-driven turbulence. MRI requires sufficient ionisation for strong gas–field coupling. The geometric set-up adopted here is sketched on the left part of Fig. 1.
@@
The velocity amplitude depends both on $\alpha_{\rm eff}$ and on how this efficiency partitions between spatial and velocity scales, as was discussed in Hughes et al. (2011). If eddies act on scales of $\sim H$, one expects $varv_{\rm turb}\sim\alpha_{\rm eff}c_{s}$; if $\alpha_{\rm eff}$ distributes similarly across length and velocity, then $varv_{\rm turb}\sim\sqrt{\alpha_{\rm eff}}c_{s}$. Empirically, observations and simulations rather support
@@
We evaluated $\alpha_{\rm eff}$ using the non-ideal, MRI-based prescription of Thi et al. (2019), described in Appendix A. This yields $varv_{\rm turb}\approx 0.1\text{–}0.4c_{s}$ in the upper layers and $varv_{\rm turb}\lesssim 0.01c_{s}$ near the mid-plane, consistent with observational constraints, and provides the values of the $L$ parameter encoding the reconnecting current sheet length and the turbulent speed used below in section 2.3.2 to connect turbulence to reconnection geometry and, ultimately, to the non-thermal particle distribution.
@@ 2.3.1 Magnetohydrodynamic turbulence and magnetic reconnection
Turbulence by itself does not guarantee magnetic reconnection. However, the atmospheric layers of protoplanetary discs combine physical conditions that make reconnection a natural outcome of the dynamics. Indeed, in this region, irradiation by stellar far UV and X-rays maintains high ionisation fractions ($\gtrsim 0.1$, see the H+/H transition Fig. 1) ensuring tight magnetic coupling (weak Ohmic and ambipolar slippage). Where coupling is good, the MRI efficiently sustains turbulence and Maxwell stresses. The canonical expectation is that MRI is most efficient for high $\beta_{p}$ (¿100); MRI in low-$\beta_{p}$ require additional conditions to grow. The parameter $\beta_{p}=P_{\text{gas}}/P_{\text{mag}}$ is the ratio of the gas pressure to the magnetic pressure. In particular, Kim and Ostriker (2000) showed that reducing the toroidal component of the magnetic field re-opens the MRI window even at low $\beta_{p}$. Recent simulations of disc surface layers are consistent with this picture. Under conditions of strong ionisation and suitable field geometry, MRI-driven turbulence naturally generates magnetic reversals, shears and then thin current sheets, the precursors of reconnection (e.g. Jacquemin-Ide et al., 2021; Rosenberg and Ebrahimi, 2021), making this process frequent and energetically relevant in disc atmospheres.
@@
There are only few works linking turbulence and magnetic reconnection in accretion discs. Kadowaki et al. (2018) explore the connection between turbulence and fast magnetic reconnection in accreting systems. The authors conduct 3D ideal MHD simulations in the shearing box approximation. Although ideal, the simulations allow magnetic reconnection thanks to numerical resistivity. Their set-up starts from an initial azimuthal magnetic field and an exponential profile of the gas density over a scale $H$. The azimuthal magnetic field component dominates the initial radial and vertical components. The initial plasma beta parameter, $\beta_{p}$, is set to 1, 10 and 100 in a series of three different runs. The simulations follow the evolution of the magnetic components under the effect of the MRI but also of the Parker-Rayleigh-Taylor instability (PRTI, Parker, 1966).
@@
The PRTI together with MRI, sustains an $\alpha-\Omega$ dynamo that reaches a quasi-steady state in which the characteristic reconnection speed is $V_{r}\approx(0.13\pm 0.09)V_{A}$ in the disc and $(0.17\pm 0.10)V_{A}$ in the corona, and the typical current-sheet width is $\Delta_{r}\approx 0.08H$ (disc) and 0.10H0.10\,H (corona), where $V_{A}$ is the local Alfvén speed. We computed the scale height using the Keplerian thin disc relation $H=c_{s}/\Omega_{K}$.  [^3] Although the ProDiMo structure is moderately flared ($H/R\!\sim\!0.1$), the deviation of the angular frequency from Keplerian due to pressure support is $\mathcal{O}[(H/R)^{2}]\!\sim\!10^{-2}$ (Armitage, 2019). This percent-level correction is well within typical astrophysical uncertainties, so the Keplerian approximation is adequate for our purposes.
@@
The values of $V_{r},\Delta_{r},H,V_{A},M_{A}$} computed from ProDiMo thus provide a coherent, simulation-supported model from which we estimate the reconnection–driven EP acceleration parameters, as is detailed in Sect. 3.2.
@@ 2.3.2 Multi-scale magnetic reconnection modelling and particle acceleration
At scales larger than kinetic ones, magnetic reconnection is allowed through the breaking of magnetic frozen-in conditions by collisions, even when the plasma is fully ionised and neutrals are not present (see e.g. Sweet 1958; Furth et al. 1963). While at kinetic scales magnetic reconnection is known to be sustained by electron inertia or the electron pressure tensor and to be efficient (i.e. fast) in converting magnetic energy into particle acceleration, fluid models have struggled to provide an explanation for fast energy release. Different models have been proposed in recent years (see e.g. the discussion in Lazarian et al. 2020) , of which the tearing instability on thin current sheets is the most promising (see e.g. Loureiro et al. 2007; Bhattacharjee et al. 2009; Pucci and Velli 2014). The survival of laminar current sheets, necessary for the onset of the tearing instability, remains controversial in the context of turbulent protoplanetary discs, and outside the scope of this work. Still, reconnection is observed in 2D and also 3D turbulent simulations even with sufficiently high magnetic Reynolds number (see Dong et al. 2022 and references therein). In the latter work, the formation of reconnection-produced magnetic flux ropes in 3D is shown to exhibit a complex morphology, differing significantly from the simpler island-like morphology in 2D. The authors note the similar size of the flux ropes between the 2D and 3D configurations, though the size of plasmoids is distributed over a wide variety of scales.
@@
The length, $L_{r}$, of the largest current sheet forming in a turbulent region is assumed to be approximately equal to the local turbulent injection scale, i.e. $L\approx L_{r}$; the largest current sheets would be destroyed by the turbulence. The thickness, $\Delta_{r}$, is set by the scale over which a magnetic field line deviates from its original direction (see Lazarian and Vishniac 1999) or, adopting a laminar tearing model, the intrinsic scale at which reconnection becomes fast (Bhattacharjee et al., 2009; Loureiro et al., 2007; Pucci and Velli, 2014). Fig. 2 schematically illustrates our working picture: sub-Alfvénic MRI turbulence ($M_{A}<1$) at the injection scale, $L$ (left), generates current sheets where turbulent magnetic reconnection operates (right). The reconnecting current sheet has a typical length, $L_{r}$, and thickness, $\Delta_{r}$, with $L_{r}=L$.
@@
Figure 2: Sketch of MRI-driven turbulence and embedded turbulent magnetic reconnection. Left: Eddies at the injection scale $L$ in a sub-Alfvénic regime characterised by the Alfvénic Mach number $M_{A}$. Right: Reconnection layer produced by the turbulence, with length $L_{r}=L$ and thickness $\Delta_{r}$. The turbulence parameters $L,M_{A}$} set the reconnection, geometry parameters $L_{r},\Delta_{r}$}, speed $V_{\rm rec}\sim V_{A}M_{A}^{2}$, and EP acceleration parameters used in this work.
@@
In a generic magnetic reconnection geometry, EP acceleration is characterised by three parameters – the acceleration rate $\alpha_{\rm acc}$, the escape time, $\tau_{\rm esc}$, and the injection time, $\tau_{\rm inj}$ – which depend on the reconnection model and on the characteristic size and lifetime of the reconnecting current sheet (Lazarian and Vishniac, 1999; Kowal et al., 2012; Lazarian et al., 2020; Xu and Lazarian, 2023),
@@
For the purpose of this paper we consider the developments of the turbulent-reconnection framework of Zhang et al. (2023) to be appropriate for the estimation of the reconnection parameters. Thus, the turbulent parameters $L,M_{A}$} set the reconnection speed, $V_{\rm rec}\sim V_{A}M_{A}^{2}$ (Lazarian and Vishniac, 1999), and, as is shown below, the parameters of the EP acceleration model. The particles are considered to be accelerated multiple times within the turbulent layer, transitioning between different current sheets, and undergo Fermi acceleration through island compression.
The same assumption holds when assuming that the plasmoid instability or any fast tearing instability kicks in. In particular we consider here the ’ideal’ tearing instability (see Pucci and Velli 2014). Then the acceleration rate is $\alpha_{\rm acc}\sim t^{-1}_{A}$, where the proportionality constant depends on the topology of the reconnecting magnetic field (Pucci et al., 2018) and other non-ideal effects (see e.g. Pucci et al. 2017; Shi et al. 2019). In this work, we adopt the results of Xu and Lazarian (2023), giving
@@
In conclusion, comparing collisional fluid modelling to Zhang et al. (2023) brings correction factors to the energy effectively converted within the turbulent layer in the disc’s atmosphere. The development of detailed particle acceleration models based on the interplay of reconnection and turbulence remains largely phenomenological, and various types of dimensional analysis may be adopted. Not knowing the details of the microscopic acceleration process at this stage, we introduce the efficiency parameter, $\kappa$, and defer a full treatment to future work.
@@ 3.1 Energy distribution from acceleration by a Fermi-like process
Magnetic reconnection can efficiently accelerate a fraction of the thermal particle population to supra-thermal energies through Fermi-like acceleration mechanism. Assuming such an acceleration process is active in the region of interest, the energy gain by a particle is proportional to the particle kinetic energy, $\dot{E}=\alpha_{\rm acc}E$, with $\alpha_{\rm acc}$ being the acceleration rate. The equation for the energy distribution is (Guo et al., 2014, 2015; Oka et al., 2023)
@@
where $F_{\rm nt}$ is the particle energy distribution (erg-1 cm-3), $\dot{E}$ is the particle energy gain rate (erg s-1), $\tau_{\rm inj}$ is the timescale for injection of the initial particle distribution, $F_{\rm inj}$, into the accelerating region, and $\tau_{\rm esc}$ is the escape timescale from the acceleration region.
Assuming that the initial upstream distribution is a Maxwellian distribution  [^4] The analysis performed in Guo et al. (2014) has been developed in the context of plasmas at high magnetisation, but the general form of the solution does not depend on the magnetisation regime, it simply assumes Fermi-I type acceleration, an injection and an escape rate. The solution in Eq. 8 assumes non-relativistic temperature of the background plasma, which is the case in the astrophysical system under consideration.,
@@
where $F_{\rm inj}$ is the upstream particle energy distribution in units of thermal energy, $\epsilon$ is the energy in units of thermal energy, $\epsilon=E/E_{\rm th}$, with $E_{\rm th}=3/2k_{B}T$, where $k_{B}$ is the Boltzmann constant and $T$ the temperature of the medium, and $n_{\rm inj}$ is the density of particles entering in the acceleration region.
The solution of Eq. (6) is a non-thermal distribution that can be written in units of thermal energy as
@@
In the next sections, we evaluate the acceleration parameters assuming a specific turbulent magnetic reconnection model and see how to normalise it based on the power released by the turbulence in the disc.
@@ 3.2 Energy distribution parameters in turbulent magnetic reconnection
The shape of the non-thermal energy distribution is set by the three parameters, $\alpha_{\rm acc}$, $\tau_{\rm esc}$, and $\tau_{\rm inj}$, as defined in Eq. (5), which depend on the characteristic size and lifetime of the reconnecting current sheet.
@@
After a time, $\tau_{\rm CS}$, a power-law component emerges if the factor $\Gamma_{3/2+\beta}\left(\epsilon e^{-\alpha_{\rm acc}\tau_{\rm CS}}\right)-\Gamma_{3/2+\beta}(\epsilon)$ in Eq. (8) is (approximately) energy-independent. This holds across the supra-thermal range provided $\epsilon e^{-\alpha_{\rm acc}\tau_{\rm CS}}\lesssim 1$. For $\epsilon e^{-\alpha_{\rm acc}\tau_{\rm CS}}\gtrsim 1$, the spectrum bends away from a pure power law, defining a finite-time cut-off,
@@
which directly reflects Fermi-like energisation, with $\dot{\epsilon}=\alpha_{\rm acc}\epsilon$ acting for a duration, $\tau_{\rm CS}$.
@@
At the low-energy end, we define the injection threshold, $\epsilon_{\rm inj}$, by continuity between thermal and non-thermal components,
@@
With the spectral shape fixed by $\alpha_{\rm acc}$, $\tau_{\rm inj}$, and $\tau_{\rm esc}$ (thus setting $p$, $\epsilon_{\rm inj}$, and $\epsilon_{\max}$), the remaining task is to determine the normalisation, i.e. the number density of non-thermal particles, from the local energy budget. In the next subsection, we link the particle power to a fraction of the turbulent accretion dissipation to obtain the non-thermal EP density.